# A simple model of globally magnetized accretion discs

### v1

@@ 7.2 Global considerations and time-dependence
Our analysis strongly constrains the disc behaviors compatible with specified outer boundary conditions. Under highly magnetized conditions and simple assumptions about disc turbulence, the accretion rate is a function of the magnetic field components, $\dot{M}\sim rB_{\phi}B_{z}/\Omega$, where $B_{z}$ is imposed and $B_{\phi}$ is determined by dynamo action. For “thinnish” discs with $H/r\la$ a few tenths, we argue in §3 that $B_{\phi}\sim 10B_{z}$, leading to the estimate of $\dot{M}_{\rm mag}$ in equation (62). This is the maximum rate at which accretion can be driven by large-scale magnetic torques. Thus, if an accretion flow is set up with a specified $B_{0}$ and $\dot{M}_{0}$ at radius $r_{0}$, and $\dot{M}_{0}$ exceeds the maximum magnetically boosted rate, then steady-state accretion can only occur through internal viscous stresses, i.e., as a Shakura-Sunyaev disc strongly dominated by gas pressure ($\beta_{z}\gg 1$), with
@@ Acknowledgements
I thank Phil Armitage, Jason Dexter, Geoffroy Lesur, Chris Reynolds, and Nico Scepi for valuable discussions during the preparation of this paper, and acknowledge financial support from NASA Astrophysics Theory Program grants NNX17AK55G and 80NSSC22K0826, and NSF Grant AST-1903335.

### v2

@@ 6.2 Global considerations and time-dependence
Our analysis strongly constrains the disc behaviors compatible with specified outer boundary conditions. Under highly magnetized conditions and simple assumptions about disc turbulence, the accretion rate is a function of the magnetic field components, $\dot{M}\sim rB_{\phi}B_{z}/\Omega$, where $B_{z}$ is imposed and $B_{\phi}$ is determined by the balance between shear-amplification and vertical diffusion. For “thinnish” discs with $H/r\la$ a few tenths, we argue in §3 that $B_{\phi}\sim 10B_{z}$, leading to the estimate of $\dot{M}_{\rm mag}$ in equation (54). This is the maximum rate at which accretion can be driven by large-scale magnetic torques. Thus, if an accretion flow is set up with a specified $B_{0}$ and $\dot{M}_{0}$ at radius $r_{0}$, and $\dot{M}_{0}$ exceeds the maximum magnetically boosted rate, then steady-state accretion can only occur through internal viscous stresses, i.e., as a Shakura-Sunyaev disc strongly dominated by gas pressure ($\beta_{z}\gg 1$), with
@@ Acknowledgements
I thank Phil Armitage, Jason Dexter, Geoffroy Lesur, Chris Reynolds, and Nico Scepi for valuable discussions during the preparation of this paper, and the anonymous referee for thoughtful and constructive criticism. I acknowledge financial support from NASA Astrophysics Theory Program grants NNX17AK55G and 80NSSC22K0826, and NSF Grant AST-1903335.
@@
Appendix A Stability
While a full stability analysis of our disc evolution equations is beyond the scope of this paper, we can examine the quasi-local stability of our steady-state models (§4) by perturbing them with short-wavelength radial perturbations of the form $y=y_{0}(1+\delta y)$ with $\delta y\propto e^{i(kr+\omega t)}$ and $H\ll k^{-1}\ll r$. The lower limit on $k^{-1}$ is intended to separate the scale of the perturbations from the presumed scale of velocity and magnetic fluctuations responsible for turbulent diffusivity. We take $\delta\xi=0$, which is plausible since the external field configuration (and hence the inclination of the poloidal field at the disc surface) is probably governed by global external boundary conditions. Defining $\nu\equiv(\omega/k)(H_{0}/\eta_{0})$, we write the perturbed mass and flux equations in the form
$\nu(\delta\rho+\delta H)=\delta\rho+\delta H+\delta\bar{v}\ ,$ (59)
$\nu\delta B_{z}=-\delta v_{B}\ ,$ (60)
$\nu\left(\delta v_{{\rm A}z}+{1\over 2}\delta\rho\right)=\left({\pi\over 2}\right)^{2}\delta q\ .$ (63)
A.1 Weakly magnetized discs
A.2 Strongly magnetized discs
A self-consistency condition on this stable regime leads to a lower limit on disc thickness. For $c_{s}\la v_{{\rm A}\phi}\la(c_{s}v_{k})^{1/2}$, the maximum MRI growth rate decreases with $v_{{\rm A}\phi}$ (Blaes & Balbus, 1994), actually vanishing for $v_{{\rm A}\phi}=2\sqrt{2}(c_{s}v_{K})^{1/2}$ when $B_{\phi}\propto r^{-5/4}$ (Pessah & Psaltis, 2005; Das et al., 2018). According to the saturation argument proposed by Begelman & Armitage (2023), $\eta\propto v_{{\rm A}\phi}$ in this regime (provided that MRI is responsible for the turbulence), implying $b\sim-1$ and instability. In order to enter the stable regime of disc evolution, $H$ must satisfy